# Generalized finite and affine $W$-algebras in type $A$

### v1

@@ 1 Introduction
In the context of the vertex algebra theory, affine $W$-algebras were used in [2] to describe the centers of the universal affine vertex algebras at the critical level associated with the centralizers $\mathfrak{a}$. This description was further applied to quantize the Mishchenko–Fomenko subalgebras in $U(\mathfrak{a})$. The center at the critical level in type $A$ was further investigated in [13], where its explicit generators were produced. It was shown in [14] that analogues of the affine $W$-algebras can be associated with the underlying Lie algebras $\mathfrak{a}$ and their description in terms of generators was also given.
@@
Our goal in this paper is to apply a version of the BRST complex $C^{k}(\lambda,\mu)$ of the quantum Drinfeld–Sokolov reduction to construct a new family of affine $W$-algebras $W^{k}(\lambda,\mu)$ associated with the centralizers $\mathfrak{a}$ of nilpotent elements in the Lie algebra $\mathfrak{gl}_{N}$. Here $\lambda$ is a partition of $N$ with $n$ parts corresponding to the chosen nilpotent element, while $\mu$ is a partition of $n$. We show that $W^{k}(\lambda,\mu)$ inherits a vertex algebra structure from $C^{k}(\lambda,\mu)$ (Theorem 3.5). The main results concerning the structure of $W^{k}(\lambda,\mu)$ are given in Theorem 3.10 and Corollary 3.11, where its generating sets were described.
@@
The family of affine $W$-algebras $W^{k}(\lambda,\mu)$ turns out to interpolate between several classes of vertex algebras previously studied in the literature. In the specialization where $\lambda=(1^{N})$ is the column-partition of $N$ and $\mu=(N)$ is the row-partition, we get the affine $W$-algebra $W^{\hskip 1.0ptk}(\mathfrak{gl}_{N})$ going back to the pioneering works of Zamolodchikov [20] and Fateev and Lukyanov [7]. Such $W$-algebras associated with arbitrary simple Lie algebras originate in conformal field theory and they were intensively studied both in mathematics and physics literature; see e.g. [1], [3] and [8, Ch. 15] for detailed expositions and reviews.
Furthermore, if $\lambda$ is the column-partition while the partition $\mu$ of $N$ is arbitrary, then $W^{k}(\lambda,\mu)$ coincides with the affine $W$-algebra $W^{k}(\mathfrak{gl}_{N},f)$ associated with a nilpotent element $f\in\mathfrak{g}$ of type $\mu$, as introduced by Kac, Roan and Wakimoto [11]; see Remark 3.12 below.
@@
In a different specialization where the partition $\lambda$ with $n$ parts is arbitrary and $\mu=(n)$ is the row-partition, $W^{k}(\lambda,(n))$ is the affine $W$-algebra $W^{k}(\mathfrak{a})$ introduced in a recent work of the second author [14]. For an alternative choice, where $\mu=(1^{n})$ is the column-partition, the $W$-algebra $W^{k}(\lambda,(1^{n}))$ coincides with the universal affine vertex algebra $V^{k}(\mathfrak{a})$, where $\mathfrak{a}$ is the centralizer of a nilpotent element of type $\lambda$. Note that, in general, the $W$-algebra $W^{k}(\lambda,\mu)$ need not be conformal; see Example 5.2.
@@
Using the same partitions $\lambda$ and $\mu$ as above, we also introduce a family of associative algebras $U(\lambda,\mu)$ which we call the generalized finite $W$-algebras; see Definition 2.1. Our main result concerning these objects is Theorem 4.6 which connects the affine $W$-algebra $W^{k}(\lambda,\mu)$ with $U(\lambda,\mu)$ via the Zhu functor originated in [21]. It was already established by De Sole and Kac [6], that the application of the Zhu functor to the affine $W$-algebra $W^{\hskip 1.0ptk}(\mathfrak{gl}_{N},f)$ associated with a nilpotent element $f\in\mathfrak{gl}_{N}$ of type $\mu$ yields the finite $W$-algebra whose definition goes back to Kostant [12]; see [9] and [17]. In our notation, it coincides with $U(\lambda,\mu)$ with $\lambda=(1^{N})$; see Example 2.2 below. Our Theorem 4.6 dealing with arbitrary $\lambda$ and $\mu$ thus provides a generalization of this property and is based on arguments similar to [6]. Note that the finite $W$-algebras $U(\lambda,\mu)$ with $\lambda=(1^{N})$ were also described from a different viewpoint in [5] (denoted there by $W(\pi)$ for the pyramid $\pi$ associated with $\mu$) by using the shifted Yangians; the particular case of rectangular pyramids appeared earlier in [18].
@@
As with the affine case, the algebras $U(\lambda,\mu)$ interpolate between several well-studied objects. The specialization with $\mu=(1^{n})$ and arbitrary $\lambda$ recovers the universal enveloping algebra ${\rm U}(\mathfrak{a})$ of the centralizer $\mathfrak{a}$ of a nilpotent element of type $\lambda$ (Example 2.3). The specialization where $\mu=(n)$ (which we refer to as the ‘principal nilpotent case’) is considered in Theorem 5.4, where we obtain an explicit description of generators of $U(\lambda,(n))$, then show that this algebra is isomorphic to the center of ${\rm U}(\mathfrak{a})$ (Corollary 5.6).
The very last Section 5.2 is devoted to the ‘minimal nilpotent case’ with $\mu=(1^{n-2}2)$ and arbitrary $\lambda$. We give explicit descriptions of generators of both the finite and affine $W$-algebras $U(\lambda,\mu)$ and $W^{k}(\lambda,\mu)$; see the respective Theorems 5.7 and 5.8.
@@ 2 Generalized finite $W$-algebras
$e=\sum_{\begin{subarray}{c}i=1,\dots,N-1\\[1.39998pt] \text{row}_{\lambda}(i)=\text{row}_{\lambda}(i+1)\end{subarray}}e_{i\,i+1},$ (2.2)
@@
$E_{ij}^{(r)}=\sum_{\begin{subarray}{c}\text{row}_{\lambda}(a)=i,\,\text{row}_{\lambda}(b)=j\\[1.39998pt] \text{col}_{\lambda}(b)-\text{col}_{\lambda}(a)=r\end{subarray}}e_{ab}.$ (2.5)
@@
$\mathsf{e}:=\sum_{\begin{subarray}{c}i=1,\dots,n-1\\[1.39998pt] \text{row}_{\mu}(i)=\text{row}_{\mu}(i+1)\end{subarray}}E_{i\,i+1}^{(\lambda_{i+1}-1)}\in\mathfrak{a}(1)$ (2.9)
@@
$\mathsf{e}=0,\qquad\mathsf{e}=E_{23}^{(4)}\qquad\text{and}\qquad\mathsf{e}=E_{12}^{(2)}+E_{23}^{(4)}.$
@@
$\chi\big{(}E_{i\,i+1}^{(\lambda_{i+1}-1)})=1$ (2.10)
if $\text{row}_{\mu}(i)=\text{row}_{\mu}(i+1)$ with $i\in\{1,\dots,n-1\}$, and $\chi(E_{ij}^{(r)})=0$ for all remaining triples $(i,j,r)\in S_{\lambda,\mu}$. Consider the left ideal
@@
of the universal enveloping algebra $U(\mathfrak{a})$. It is clear that the adjoint action of $\mathfrak{n}_{\lambda,\mu}$ on $U(\mathfrak{a})$ induces the action of $\mathfrak{n}_{\lambda,\mu}$ on the quotient $U(\mathfrak{a})/\mathcal{I}_{\lambda,\mu}$. As with the finite $W$-algebras (cf. [5] and [6, Appendix]), the subspace of $\mathfrak{n}_{\lambda,\mu}$-invariants turns out to be an associative algebra with the product inherited from $U(\mathfrak{a})$ as we verify below.
@@
To verify that $U(\lambda,\mu)$ is a well-defined associative algebra, take $A,B\in U(\mathfrak{a})$ such that $\overline{A},\overline{B}\in U(\lambda,\mu)$ are their images in the quotient, and $\mathsf{n}\in\mathfrak{n}_{\lambda,\mu}$. It is sufficient to show that $[\mathsf{n},AB]\in\mathcal{I}_{\lambda,\mu}$. Since $\overline{A}\in U(\lambda,\mu)$, we have $[\mathsf{n},A]=a(\mathsf{m}+\chi(\mathsf{m}))$ for some $a\in U(\mathfrak{a})$ and $\mathsf{m}\in\mathfrak{n}_{\lambda,\mu}$. By the Leibniz rule,
@@
Let $\lambda=(1,1,\dots,1)$ be the column-pyramid with $N$ boxes so that $\mathfrak{a}=\mathfrak{gl}_{N}$ and let $\mu$ be a pyramid with $N$ boxes. Consider the element $\mathsf{e}$ associated with $\mu$ as defined in (2.9). Then
$U(\lambda,\mu)=U(\mathfrak{g},\mathsf{e})$
is the finite $W$-algebra associated with $\mathfrak{g}$ and $\mathsf{e}$; see [9], [12] and [17].

### v2

@@ 1 Introduction
1.1 $W$-algebras
The appearance of affine $W$-algebras was motivated by physics, originating in the work of Zamolodchikov [31]; he extended the Virasoro algebra by adding fields of higher conformal dimension thus discovering the $W_{3}$ algebra associated with the Lie algebra $\mathfrak{sl}_{3}$. The $W_{N}$-algebras corresponding to $\mathfrak{sl}_{N}$ were then constructed by Fateev and Lukyanov [12]. Due to their significance in the conformal field theory, the $W$-algebra symmetries were since extensively studied in the physics literature; see e.g. [5] for a review.
A conceptual mathematical definition of the $W$-algebras in the context of the vertex algebra theory is due to Feigin and Frenkel [13] who introduced them via the quantized Drinfeld–Sokolov reduction. According to this definition, the $W$-algebra $W^{k}(\mathfrak{g})$ at the level $k\in\mathbb{C}\hskip 1.0pt$ corresponding to the simple Lie algebra $\mathfrak{g}$ is the zeroth cohomology of the BRST complex of the quantum Drinfeld–Sokolov reduction; see [3] and [15, Ch. 15] for detailed expositions.
The theory was further developed by Kac, Roan and Wakimoto [20] introducing a wider class of $W$-algebras $W^{k}(\mathfrak{g},f)$ associated with simple Lie algebras and superalgebras $\mathfrak{g}$ and nilpotent elements $f\in\mathfrak{g}$. For a principal nilpotent $f$ the algebra $W^{k}(\mathfrak{g},f)$ coincides with $W^{k}(\mathfrak{g})$ and for the trivial nilpotent $f$ the corresponding $W$-algebra $W^{k}(\mathfrak{g},f)$ is the affine vertex algebra $V^{k}(\mathfrak{g}).$ In other words, for given Lie superalgebra $\mathfrak{g}$, the family of $W$-algebras interpolate the principal $W$-algebra and affine vertex algebra. In the principal case, the investigation of the structural theory and representations of $W$-algebras led Frenkel, Kac and Wakimoto [16] to remarkable conjectures concerning their rationality and the existence and description of modular invariant representations. The $W$-algebra representation theory was further developed by Arakawa in [1] and [2], where both conjectures were proved.
In the limit $k\to\infty$ the algebra $W^{k}(\mathfrak{g})$ becomes the classical $W$-algebra $W(\mathfrak{g})$ which is a commutative algebra equipped with a Poisson bracket. Its definition goes back to Drinfeld and Sokolov [11]; they used the algebras $W(\mathfrak{g})$ to introduce equations of the KdV type for arbitrary simple Lie algebras. In a more recent work by De Sole, Kac and Valeri [10] the construction of Drinfeld and Sokolov (generalized to an arbitrary nilpotent element $f$) was described in the framework of Poisson vertex algebras. This description was then applied to construct integrable hierarchies of bi-Hamiltonian equations.
According to the Feigin–Frenkel duality property [13], another way to recover the classical $W$-algebra $W({}^{L}\mathfrak{g})$ associated with the Langlands dual Lie algebra $^{L}\mathfrak{g}$ is to take the critical level for $W^{k}(\mathfrak{g})$; i.e., to evaluate $k$ at the negative of the dual Coxeter number of $\mathfrak{g}$. Due to the celebrated theorem of Feigin and Frenkel [14], the algebra $W({}^{L}\mathfrak{g})$ is isomorphic to the center of the affine vertex algebra at the critical level corresponding to $\mathfrak{g}$.
An independent theory of finite $W$-algebras was conceived in the work of Kostant [22]. More recent approaches were developed by Premet [27] and Gan and Ginzburg [17]. The connection between the affine and finite $W$-algebras was established by De Sole and Kac [9], and (in the principal nilpotent case) by Arakawa [1]. They showed that the application of the Zhu functor originated in [32] to the affine $W$-algebra $W^{\hskip 1.0ptk}(\mathfrak{g},f)$ associated with a nilpotent element $f\in\mathfrak{g}$ yields the finite $W$-algebra corresponding to $\mathfrak{g}$ and $f$.
1.2 Generalized $W$-algebras
@@
The main motivation of our work comes from the results of Arakawa and Premet [4]. They found a remarkable way to use affine $W$-algebras to describe the centers of the universal affine vertex algebras at the critical level associated with the centralizers $\mathfrak{a}$, thus obtaining a version of the Feigin–Frenkel theorem for centralizers. This description was further applied to quantize the Mishchenko–Fomenko subalgebras in $U(\mathfrak{a})$. The center at the critical level in type $A$ was further investigated in [23], where its explicit generators were produced. It was shown in [24] that analogues of the affine $W$-algebras can be associated with the underlying Lie algebras $\mathfrak{a}$ and their description in terms of generators was also given.
@@
Our goal in this paper is to apply a version of the BRST complex $C^{k}(\lambda,\mu)$ of the quantum Drinfeld–Sokolov reduction to construct a new family of vertex algebras $W^{k}(\lambda,\mu)$ associated with the centralizers $\mathfrak{a}$ of nilpotent elements in the Lie algebra $\mathfrak{gl}_{N}$. Here $\lambda$ is a partition of $N$ with $n$ parts corresponding to the chosen nilpotent element, while $\mu$ is a partition of $n$. We show that $W^{k}(\lambda,\mu)$ inherits a vertex algebra structure from $C^{k}(\lambda,\mu)$ (Theorem 3.6). The main results concerning the structure of $W^{k}(\lambda,\mu)$ are given in Theorem 3.12 and Corollary 3.13, where its generating sets were described.
@@
The family of affine $W$-algebras $W^{k}(\lambda,\mu)$ turns out to interpolate between several classes of vertex algebras previously studied in the literature. In the specialization where $\lambda=(1^{N})$ is the column-partition of $N$ and $\mu=(N)$ is the row-partition, we get the affine $W$-algebra $W^{\hskip 1.0ptk}(\mathfrak{gl}_{N})$ going back to [31] and [12]. If the partition $\mu$ of $N$ is arbitrary, then $W^{k}(\lambda,\mu)$ coincides with the affine $W$-algebra $W^{k}(\mathfrak{gl}_{N},f)$ associated with a nilpotent element $f\in\mathfrak{gl}_{N}$ of type $\mu$, as introduced in [20]; see Remark 3.14 below.
@@
In a different specialization where the partition $\lambda$ with $n$ parts is arbitrary and $\mu=(n)$ is the row-partition, $W^{k}(\lambda,(n))$ is the affine $W$-algebra $W^{k}(\mathfrak{a})$ introduced in [24]. For an alternative choice, where $\mu=(1^{n})$ is the column-partition, the $W$-algebra $W^{k}(\lambda,(1^{n}))$ coincides with the universal affine vertex algebra $V^{k}(\mathfrak{a})$, where $\mathfrak{a}$ is the centralizer of a nilpotent element of type $\lambda$. Note that, in general, the $W$-algebra $W^{k}(\lambda,\mu)$ need not be conformal; see Example 5.2.
@@
Using the same partitions $\lambda$ and $\mu$ as above, we also introduce a family of associative algebras $U(\lambda,\mu)$ which we call the generalized finite $W$-algebras; see Definition 2.1. Our main result concerning these objects is Theorem 4.6 which connects the affine $W$-algebra $W^{k}(\lambda,\mu)$ with $U(\lambda,\mu)$ via the Zhu functor by analogy with [9]. In our notation, the construction in loc. cit. corresponds to $\lambda=(1^{N})$ and a nilpotent element $f\in\mathfrak{gl}_{N}$ of type $\mu$; see Example 2.2 below. Note that the finite $W$-algebras $U(\lambda,\mu)$ with $\lambda=(1^{N})$ were also described from a different viewpoint in [7] (denoted there by $W(\pi)$ for the pyramid $\pi$ associated with $\mu$) by using the shifted Yangians; the particular case of rectangular pyramids appeared earlier in [28].
@@
As with the affine case, the algebras $U(\lambda,\mu)$ interpolate between several well-studied objects. The specialization with $\mu=(1^{n})$ and arbitrary $\lambda$ recovers the universal enveloping algebra $U(\mathfrak{a})$ of the centralizer $\mathfrak{a}$ of a nilpotent element of type $\lambda$ (Example 2.3). The specialization $\lambda_{1}=\lambda_{2}=\cdots=\lambda_{n}=p$ with arbitrary $\mu$ recovers the finite $W$-algebras associated with truncated current Lie algebras, also known as generalized Takiff algebras [29]; see Remark 2.4. In particular, the Whittaker modules for $U(\lambda,\mu)$ in this setting were studied in [18]. The specialization where $\mu=(n)$ (which we refer to as the ‘principal nilpotent case’) is considered in Theorem 5.4, where we obtain an explicit description of generators of $U(\lambda,(n))$, then show that this algebra is isomorphic to the center of $U(\mathfrak{a})$ (Corollary 5.6). The very last Section 5.2 is devoted to the ‘minimal nilpotent case’ with $\mu=(1^{n-2}2)$ and arbitrary $\lambda$. We give explicit descriptions of generators of both the finite and affine $W$-algebras $U(\lambda,\mu)$ and $W^{k}(\lambda,\mu)$; see the respective Theorems 5.7 and 5.8.
1.3 Further directions
As written above, in this paper, we introduce new family of vertex algebras which interpolate known vertex algebras such as $V^{k}(\mathfrak{gl}_{N})$, $V^{k}(\mathfrak{a})$, and $W^{k}(\mathfrak{gl}_{N},f).$ Furthermore, we showed some of generalized $W$-algebras are chiralization of some known associative algebras. As an example, a finite $W$-algebra associated with a Takiff algebra is obtained as a Zhu algebra of $W^{k}(\lambda,\mu)$ for a rectangular shape partition $\lambda$.
As a future project, we will observe the relation between the $W$-algebras associated with $\mathfrak{gl}_{N}$ and $\mathfrak{a}.$ In particular, when $f$ is a nilpotent element of $\mathfrak{gl}_{N}$ inducing the grading (2.7), the BRST complex for $W^{k}(\lambda,\mu)$ is contained in the BRST complex for $W^{k}(\mathfrak{gl}_{N},f)$. Moreover, the conformal vector on the $W^{k}(\mathfrak{gl}_{N},f)$ induces the conformal weight decomposition of $W^{k}(\lambda,\mu)$. Hence comparing the two vertex algebras should be an interesting problem and also this observation can provide a quantum affine analogue of the relation between $S(\mathfrak{g})^{\mathfrak{g}}$ and $S(\mathfrak{a})^{\mathfrak{a}}$ established in [26]. Another interesting possible link might be observed between $W^{k}(\lambda,\mu)$ and $W^{k}(\mathfrak{gl}_{N},f_{\lambda,\mu})$ where $f_{\lambda,\mu}$ is dual to $\mathsf{e}_{\lambda,\mu}$ in (2.9). We expect that $W^{k}(\lambda,\mu)$ embeds into $W^{k}(\mathfrak{gl}_{N},f_{\lambda,\mu})$.
We also believe that both the vertex algebras $W^{k}(\lambda,\mu)$ and associative algebras $U(\lambda,\mu)$ deserve further investigation regarding their structure theory and representations, and moreover Feigin–Frenkel-type duality, classical counterparts of $W^{k}(\lambda,\mu)$ and critical level phenomena would be interesting. Another interesting question is a possible relationship of the algebras $U(\lambda,\mu)$ with suitably modified shifted Yangians of [7].
@@ 2 Generalized finite $W$-algebras
$e=\sum_{\begin{subarray}{c}i=1,\dots,N-1\\[1.63885pt] \text{row}_{\lambda}(i)=\text{row}_{\lambda}(i+1)\end{subarray}}e_{i\,i+1},$ (2.2)
@@
$E_{ij}^{(r)}=\sum_{\begin{subarray}{c}\text{row}_{\lambda}(a)=i,\,\text{row}_{\lambda}(b)=j\\[1.63885pt] \text{col}_{\lambda}(b)-\text{col}_{\lambda}(a)=r\end{subarray}}e_{ab}.$ (2.5)
@@
$\mathsf{e}_{\lambda,\mu}:=\sum_{\begin{subarray}{c}i=1,\dots,n-1\\[1.63885pt] \text{row}_{\mu}(i)=\text{row}_{\mu}(i+1)\end{subarray}}E_{i\,i+1}^{(\lambda_{i+1}-1)}\in\mathfrak{a}(1)$ (2.9)
@@
$\mathsf{e}_{\lambda,\mu}=0,\qquad\mathsf{e}_{\lambda,\mu}=E_{23}^{(4)}\qquad\text{and}\qquad\mathsf{e}_{\lambda,\mu}=E_{12}^{(2)}+E_{23}^{(4)}.$
@@
$\chi\big(E_{i\,i+1}^{(\lambda_{i+1}-1)}\big)=1$ (2.10)
if $\text{row}_{\mu}(i)=\text{row}_{\mu}(i+1)$ with $i\in\{1,\dots,n-1\}$, and $\chi(E_{ij}^{(r)})=0$ for all remaining triples $(i,j,r)\in S_{\lambda,\mu}$. For an explanation of this choice of $\chi$, see Remark 3.11. Consider the left ideal
@@
of the universal enveloping algebra $U(\mathfrak{a})$. Because $\chi\in\mathfrak{n}_{\lambda,\mu}^{*}$ with $\chi|_{\mathfrak{a}(i)}=0$ for all $i\geqslant 2$, the adjoint action of $\mathfrak{n}_{\lambda,\mu}$ on $U(\mathfrak{a})$ induces the action of $\mathfrak{n}_{\lambda,\mu}$ on the quotient $U(\mathfrak{a})/\mathcal{I}_{\lambda,\mu}$. As with the finite $W$-algebras (cf. [7] and [9, Appendix]), the subspace of $\mathfrak{n}_{\lambda,\mu}$-invariants turns out to be an associative algebra with the product inherited from $U(\mathfrak{a})$ as we verify below.
@@
To verify that $U(\lambda,\mu)$ is a well-defined associative algebra, take $A,B\in U(\mathfrak{a})$ such that $\overline{A},\overline{B}\in U(\lambda,\mu)$ are their images in the quotient, and $\mathsf{n}\in\mathfrak{n}_{\lambda,\mu}$. It is sufficient to show that $\mathsf{n},AB]\in\mathcal{I}_{\lambda,\mu}$. Since $\overline{A}\in U(\lambda,\mu)$, we have $\mathsf{n},A]=a(\mathsf{m}+\chi(\mathsf{m}))$ for some $a\in U(\mathfrak{a})$ and $\mathsf{m}\in\mathfrak{n}_{\lambda,\mu}$. By the Leibniz rule,
@@
Let $\lambda=(1,1,\dots,1)$ be the column-pyramid with $N$ boxes so that $\mathfrak{a}=\mathfrak{gl}_{N}$ and let $\mu$ be a pyramid with $N$ boxes. Consider the element $\mathsf{e}_{\lambda,\mu}$ associated with $\mu$ as defined in (2.9). Then
$U(\lambda,\mu)=U(\mathfrak{g},\mathsf{e}_{\lambda,\mu})$
is the finite $W$-algebra associated with $\mathfrak{g}$ and $\mathsf{e}_{\lambda,\mu}$; see [17], [22] and [27].
@@
Remark 2.4.
Let $N=np$ and $\lambda_{1}=\lambda_{2}=\cdots=\lambda_{n}=p$ and $\mu$ be a pyramid with $n$ boxes. We have a Lie algebra isomorphism
$\mathfrak{a}\ \cong\ \mathfrak{gl}_{n}[v]/(v^{p}),\qquad E_{ij}^{(r)}\longleftrightarrow e_{ij}v^{r}$ (2.11)
Here, $\mathfrak{gl}_{n}[v]/(v^{p})$ denotes the truncated current Lie algebra of level $p$, equipped with the Lie bracket induced from the current Lie algebra $\mathfrak{gl}_{n}[v]$. The $W$-algebras associated with $\mathfrak{gl}_{n}[v]/(v^{p})$ and a nilpotent element $e$ were constructed in [18]. The $W$-algebra in [18] is identical to the $W$-algebra constructed in this paper from the Lie algebra $\mathfrak{a}$, after identifying via (2.11). (See [8] and references therein.)